For each test statistic below, give the p value (do not round p): a. t =...

  1. For each test statistic below, give the p value (do not round p): a. t = 2.20, df = 15, one-tailed test
    b. Z = 1.10, one-tailed test
    c. t = -2.65, df = 20, two-tailed test

    d. Z = -.26, two-tailed test
Answer #1

Solution:

a)

t = 2.20, df = 15, one-tailed test

P-value = 0.021948

b)

Z = 1.10, one-tailed test

P-value = P(Z < z )

= P(Z < 1.10 )

= 0.8643

c)

t = -2.65, df = 20, two-tailed test

P-value = 0.015366

d)

Z = -.26, two-tailed test

P-value = 2 * P(Z < z )

= 2 * P(Z < -0.26 )

= 2 * 0.3974

= 0.7948